OpenStudy (anonymous):

Susan is 10 years older than Matthew.In 4 years, Susan will be twice as old as Matthew.How old are they now?

OpenStudy (slaaibak):

susan's age now = x matthews age now: x - 10 x + 4 = 2(x - 10 + 4) x + 4 = 2x - 20 + 8 -x = -16 So susan is 16 now Matthew is now 16 - 10 = 6.